# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ C.4  Convergence rate of the Doob’s guidance estimation
$\lesssim\frac{1}{\lambda}\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+\frac{1}{\lambda}\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{1}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda}{\sigma_{T-t}^{8}}.$
By setting $N=n^{\frac{1}{d+4}}$ and $\lambda=n^{-\frac{1}{d+4}}$, we have
(C.34) ${\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{2}{d+4}}\log^{2}n,$
${\mathbb{E}}\big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{1}{d+4}}\log^{2}n.$
@@
$\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{1}{d+4}}\log^{2}n,$
where the second inequality is owing to Lemma C.1, and the last inequality holds from (C.34). This completes the proof. ∎
Appendix D Error Bounds of the Controllable Diffusion Models
D.1  Error decomposition of the controllable diffusion models
@@
$\quad+\underbrace{\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{base}}^{2}}_{\text{(II)}}+\underbrace{d\exp(-T)}_{\text{(III)}}+\underbrace{\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}}_{\text{(IV)}},$
@@
According to Chen et al. [2023, Proposition C.3], we have
(D.1) $\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\underbrace{\mathrm{KL}(q_{T}\|\gamma_{d})}_{\text{initial error}}+\underbrace{\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]}_{\text{base score error}}$
@@
Step 1. Initial error in (D.1). Using Chen et al. [2023, Lemma C.4], we have
(D.2) $\mathrm{KL}(q_{T}\|\gamma_{d})\lesssim d\exp(-T).$
Step 2. Base score error and Doob’s guidance error in (D.1). Under Assumption 2, it is apparent that ${\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]\geq\underline{B}$ and $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h^{*}(t,\mathbf{x})\leq\bar{B}$ for each $t\in(T_{0},T)$. Hence, the density ratio is uniformly bounded
(D.3) $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}\frac{q_{T-t}(\mathbf{x})}{p_{T-t}(\mathbf{x})}=\frac{h^{*}(t,\mathbf{x})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]}\leq\frac{\bar{B}}{\underline{B}}.$
For the base score error term in (D.1), it follows for each $0\leq k\leq K-1$ that
@@
where the inequality holds from (D.3) and Hölder’s inequality. Consequently,
@@
(D.4) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq T\frac{\bar{B}}{\underline{B}}\varepsilon_{\mathrm{base}}^{2},$
@@
(D.5) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right].$
Step 3. Discretization error in (D.1). According to Chen et al. [2023, Lemma D.1], we have
(D.6) ${\mathbb{E}}^{{\mathbb{Q}}}\big[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\big]\lesssim\frac{dG_{k}T}{K},$
@@
$\leq\sup_{t\in(T_{0},T)}\frac{1}{\sigma_{T-t}^{2}}+\frac{\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\|\mathrm{Cov}(\mathbf{Z}_{0}|\mathbf{Z}_{t}=\mathbf{z})\|_{\mathrm{op}}$
(D.7) $\lesssim\frac{d}{\sigma_{T_{0}}^{4}},$
where the first inequality holds from the triangular inequality, and the second inequality is due to the boundedness of $\mathbf{Z}_{0}$ under Assumptions 1 and 2. Combining (D.6) and (D.7) implies
(D.8) $\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]\,\mathrm{d}t\lesssim\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}.$
Step 4. Conclusions. Substituting (D.2), (D.4), (D.5), (D.8) into (D.1) completes the proof. ∎
Corollary D.1.
@@
$\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}^{2}\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}},$
@@
$K=\Theta\Big(\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),\quad T=\Theta\Big(\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),\quad h=\frac{T}{K}=\Theta\Big(\frac{\delta^{2}}{\sigma_{T_{0}}^{4}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
$\varepsilon_{\mathrm{base}}^{2}=\Theta\Big(\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),\quad n=\Theta\Big(\frac{1}{\delta^{2(d+4)}}\log^{d+4}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
Proof of Corollary D.1.
@@
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim C^{\prime}\frac{\bar{B}T}{\underline{B}\sigma_{T_{0}}^{8}}n^{-\frac{1}{d+4}}+\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{base}}^{2}+d\exp(-T)+\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K},$
where $C^{\prime}$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. By setting
$K=\Theta\Big(\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),\quad T=\Theta\Big(\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
$\varepsilon_{\mathrm{base}}^{2}=\Theta\Big(\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),\quad n=\Theta\Big(\frac{1}{\delta^{2(d+4)}}\log^{d+4}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
@@
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}}.$
@@
D.2  Convergence rate of the controllable diffusion models
@@
(D.9) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\Big\{\sigma_{T_{0}}^{2}+\varepsilon+\frac{\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log\Big(\frac{1}{\varepsilon}\Big)\Big\},$
provided that the truncation radius $R$, the step size $h$, the number of steps $K$, the error of base score $\varepsilon_{\mathrm{base}}$, and the number of samples $n$ for Doob’s matching are set, respectively, as
$R=\Theta\Big(\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad h=\Theta\Big(\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad K=\Theta\Big(\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big)\Big),$
$\varepsilon_{\mathrm{base}}^{2}=\Theta\Big(\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad n=\Theta\Big(\frac{1}{\varepsilon^{6(d+4)}}\log^{d+4}\Big(\frac{1}{\varepsilon}\Big)\Big),$
Further, letting $T_{0}=\Theta(\varepsilon)$, it holds that
(D.10) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log\Big(\frac{1}{\varepsilon}\Big).$
@@
(D.11) ${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $=3\underbrace{{\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})}_{\text{(i)}}+3\underbrace{{\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})}_{\text{(ii)}}$
@@
Step 1. Bound the term (i) in (D.11). To estimate the 2-Wasserstein distance between the target distribution $q_{0}$ and the scaled early-stopping distribution ${\mathcal{M}}_{\sharp}q_{T_{0}}$, we begin by producing a coupling of them. Let $\mathbf{Z}_{0}\sim q_{0}$, and let ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ be independent of $\mathbf{Z}_{0}$. Define $\widetilde{\mathbf{Z}}_{T_{0}}\coloneq\mathbf{Z}_{0}+\sigma_{T_{0}}\mu_{T_{0}}^{-1}{\bm{\varepsilon}}$. It is apparent that $\widetilde{\mathbf{Z}}_{T_{0}}\sim{\mathcal{M}}_{\sharp}q_{T_{0}}$. Then
(D.12) ${\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})\leq{\mathbb{E}}\big[\|\mathbf{Z}_{0}-\widetilde{\mathbf{Z}}_{T_{0}}\|_{2}^{2}\big]=\frac{\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]=\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}.$
Step 2. Bound the term (ii) in (D.11). Let $\mathbf{Z}_{T_{0}}\sim q_{T_{0}}$. According to the definition of the truncation operator ${\mathcal{T}}_{R}$, the joint law of $(\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}},\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{Z}_{T_{0}}))$ is a coupling of ${\mathcal{M}}_{\sharp}q_{T_{0}}$ and $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$. Therefore,
@@
where the second ineq holds from Cauchy-Schwarz inequality, and the last inequality is due to Lemma E.3 and Corollary E.2. By setting $R^{2}=(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})$, we have
(D.13) ${\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})\lesssim\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon.$
Step 3. Bound the term (iii) in (D.11). Let $\mathbf{Z}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$ and $\widehat{\mathbf{Z}}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ be optimal coupled. This means
(D.14) ${\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})={\mathbb{E}}\big[\|\mathbf{Z}_{T_{0}}^{R}-\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big].$
@@
(D.15) $\leq{\mathbb{E}}\big[\|\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}^{R}-\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big]=\frac{1}{\mu_{T_{0}}^{2}}{\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}),$
where the equality holds from (D.14). Then using Villani [2009, Theorem 6.15] and the data processing inequality, we have
@@
(D.16) $\leq 2R^{2}\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}.$
Combining (D.15) and (D.16) yields

### v2

@@ D.3  Convergence rate of the Doob’s guidance estimation
$\leq\frac{C}{\lambda}\frac{\log^{4}N}{\sigma_{T-t}^{8}N^{4}}+C\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+\frac{C}{\lambda}\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{C}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{C\lambda}{\sigma_{T-t}^{8}}.$
By setting $N={\mathcal{O}}(n^{\frac{1}{d+8}})$ and $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, we have
(D.25) ${\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{4}{d+8}}\log^{4}n,$
${\mathbb{E}}\big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{2}{d+8}}\log^{4}n.$
@@
$\leq C^{\prime}\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{2}{d+8}}\log^{4}n,$
where the second inequality is owing to Lemmas C.1 and C.4, and the last inequality holds from (D.25). This completes the proof. ∎
Appendix E Derivations in Section 5.3
E.1  Error decomposition of the controllable diffusion models
@@
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]$
$\quad+\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{ref}}^{2}+d\exp(-T)+\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K},$
@@
According to Chen et al. [2023a, Proposition C.3], we have
(E.1) $\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\underbrace{\mathrm{KL}(q_{T}\|\gamma_{d})}_{\text{initial error}}+\underbrace{\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]}_{\text{base score error}}$
@@
Step 1. Initial error in (E.1). Using Chen et al. [2023a, Lemma C.4], we have
(E.2) $\mathrm{KL}(q_{T}\|\gamma_{d})\lesssim d\exp(-T).$
Step 2. Reference score error and Doob’s guidance error in (E.1). Under Assumption 2, it is apparent that ${\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]\geq\underline{B}$ and $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h^{*}(t,\mathbf{x})\leq\bar{B}$ for each $t\in(T_{0},T)$. Hence, the density ratio is uniformly bounded
(E.3) $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}\frac{q_{T-t}(\mathbf{x})}{p_{T-t}(\mathbf{x})}=\frac{h^{*}(t,\mathbf{x})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]}\leq\frac{\bar{B}}{\underline{B}}.$
For the reference score error term in (E.1), it follows for each $0\leq k\leq K-1$ that
@@
where the inequality holds from (E.3) and Hölder’s inequality. Consequently,
@@
(E.4) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq T\frac{\bar{B}}{\underline{B}}\varepsilon_{\mathrm{ref}}^{2},$
@@
(E.5) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right].$
Step 3. Discretization error in (E.1). According to Chen et al. [2023a, Lemma D.1], we have
(E.6) ${\mathbb{E}}^{{\mathbb{Q}}}\big[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\big]\lesssim\frac{dG_{k}T}{K},$
@@
(E.7) $\leq\sup_{t\in(T_{0},T)}\frac{1}{\sigma_{T-t}^{2}}+\frac{\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\|\mathrm{Cov}(\mathbf{Z}_{0}|\mathbf{Z}_{t}=\mathbf{z})\|_{\mathrm{op}}\lesssim\frac{d}{\sigma_{T_{0}}^{4}},$
where the first inequality holds from the triangular inequality, and the second inequality is due to the boundedness of $\mathbf{Z}_{0}$ under Assumptions 1 and 2. Combining (E.6) and (E.7) implies
(E.8) $\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]\,\mathrm{d}t\lesssim\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}.$
Step 4. Conclusions. Substituting (E.2), (E.4), (E.5), (E.8) into (E.1) completes the proof. ∎
Corollary E.1.
@@
$\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}^{2}\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}}\log\Big(\frac{1}{\delta}\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
@@
$T\asymp\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad K\gtrsim\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad h\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{4}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad n\gtrsim\frac{1}{\delta^{d+8}}\log^{\frac{d+8}{2}}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big).$
Proof of Corollary E.1.
@@
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\leq\frac{C}{\sigma_{T_{0}}^{8}}\Big\{\underbrace{Tn^{-\frac{2}{d+8}}\log^{4}n}_{\text{(i)}}+\underbrace{T\sigma_{T_{0}}^{8}\varepsilon_{\mathrm{ref}}^{2}}_{\text{(ii)}}+\underbrace{\sigma_{T_{0}}^{8}\exp(-T)}_{\text{(iii)}}+\underbrace{T^{2}\sigma_{T_{0}}^{4}\frac{1}{K}}_{\text{(iv)}}\Big\},$
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. By setting
$T\asymp\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad K\gtrsim\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad n\gtrsim\frac{1}{\delta^{d+8}}\log^{\frac{d+8}{2}}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),$
@@
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}}\log\Big(\frac{1}{\delta}\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big).$
@@
E.2  Convergence rate of the controllable diffusion models
@@
(E.9) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big).$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
@@
(E.10) ${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $=3\underbrace{{\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})}_{\text{(i)}}+3\underbrace{{\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})}_{\text{(ii)}}$
@@
Step 1. Bound the term (i) in (E.10). To estimate the 2-Wasserstein distance between the target distribution $q_{0}$ and the scaled early-stopping distribution ${\mathcal{M}}_{\sharp}q_{T_{0}}$, we begin by producing a coupling of them. Let $\mathbf{Z}_{0}\sim q_{0}$, and let ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ be independent of $\mathbf{Z}_{0}$. Define $\widetilde{\mathbf{Z}}_{T_{0}}\coloneq\mathbf{Z}_{0}+\sigma_{T_{0}}\mu_{T_{0}}^{-1}{\bm{\varepsilon}}$. It is apparent that $\widetilde{\mathbf{Z}}_{T_{0}}\sim{\mathcal{M}}_{\sharp}q_{T_{0}}$. Then
(E.11) ${\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})\leq{\mathbb{E}}\big[\|\mathbf{Z}_{0}-\widetilde{\mathbf{Z}}_{T_{0}}\|_{2}^{2}\big]=\frac{\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]=\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}.$
Step 2. Bound the term (ii) in (E.10). Let $\mathbf{Z}_{T_{0}}\sim q_{T_{0}}$. According to the definition of the truncation operator ${\mathcal{T}}_{R}$, the joint law of $(\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}},\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{Z}_{T_{0}}))$ is a coupling of ${\mathcal{M}}_{\sharp}q_{T_{0}}$ and $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$. Therefore,
@@
where the second ineq holds from Cauchy-Schwarz inequality, and the last inequality is due to Lemma G.3 and Corollary G.2. By setting $R^{2}=(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})$, we have
(E.12) ${\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})\lesssim\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon.$
Step 3. Bound the term (iii) in (E.10). Let $\mathbf{Z}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$ and $\widehat{\mathbf{Z}}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ be optimal coupled. This means
(E.13) ${\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})={\mathbb{E}}\big[\|\mathbf{Z}_{T_{0}}^{R}-\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big].$
@@
(E.14) $\leq{\mathbb{E}}\big[\|\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}^{R}-\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big]=\frac{1}{\mu_{T_{0}}^{2}}{\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}),$
where the equality holds from (E.13). Then using Villani [2009, Theorem 6.15] and the data processing inequality, we have
@@
(E.15) $\leq 2R^{2}\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}.$
Combining (E.14) and (E.15) yields